OpenStudy (anonymous):

5/7x-1/4x+10=9/14x these are fractions and need to solve equation please.

OpenStudy (anonymous):

x=56

OpenStudy (anonymous):

Thank You very much

OpenStudy (anonymous):

\[5/7x-1/4x+10=9/14x \] \[(13/28) x +10 = (9/14)x\] \[(13/28) x - (9/14)x = -10\] \[(-5/28)x = -10\] \[x = 280/5\] \[x=\] Your welcome.

OpenStudy (anonymous):

x=56